# Privacy-Preserving Federated Learning against Malicious Clients Based on Verifiable Functional Encryption

### v3

@@ Abstract
Federated learning is a promising distributed learning paradigm that enables collaborative model training without exposing local client data, thereby protect data privacy. However, it also brings new threats and challenges. The advancement of model inversion attacks has rendered the plaintext transmission of local models insecure, while the distributed nature of federated learning makes it particularly vulnerable to attacks raised by malicious clients. To protect data privacy and prevent malicious client attacks, this paper proposes a privacy-preserving federated learning framework based on verifiable functional encryption, without a non-colluding dual-server setup or additional trusted third-party. Specifically, we propose a novel decentralized verifiable functional encryption (DVFE) scheme that enables the verification of specific relationships over multi-dimensional ciphertexts. This scheme is formally treated, in terms of definition, security model and security proof. Furthermore, based on the proposed DVFE scheme, we design a privacy-preserving federated learning framework VFEFL that incorporates a novel robust aggregation rule to detect malicious clients, enabling the effective training of high-accuracy models under adversarial settings. Finally, we provide formal analysis and empirical evaluation of the proposed schemes. The results demonstrate that our approach achieves the desired privacy protection, robustness, verifiability and fidelity, while eliminating the reliance on non-colluding dual-server settings or trusted third parties required by existing methods.
@@ III-A System Model
Note: The superscripts of $W_{i}^{t},x_{i}^{t},D_{i}$ denote interactions and the subscripts denote client identifiers.
@@ III-C Security Model
where $\overset{c}{\equiv}$ denotes computational indistinguishability, $\lambda$ represents the security parameter, and $\hat{x}$ includes all inputs.
@@ IV-A Groups
Prime Order Group. We use a prime-order group GGen, a probabilistic polynomial-time (PPT) algorithm that, given a security parameter $\lambda$, outputs a description $\mathcal{G}=(\mathbb{G},p,g)$, where $\mathbb{G}$ is an additive cyclic group of order $p$, $p$ is a $2\lambda$-bit prime, and $g$ is the generator of $\mathbb{G}$. For any $a\in\mathbb{Z}_{p}$, given $g^{a}$, it is computationally hard to determine $a$, a problem known as the discrete logarithm problem. For any $a=\{a_{1},\cdots,a_{n}\}\in\mathbb{Z}^{n}_{p}$, $g=\{g_{1},\cdots,g_{n}\}\in\mathbb{G}^{n}$, we denote $g^{a}=\prod_{i\in[n]}g_{i}^{a_{i}}$.
@@ IV-B Zero Knowledge Proof
IV-C Verifiable Functional Encryption
Unlike traditional public key encryption, where decryption is an all-ro-nothing affair, functional encryption[FEIP] can finely controls the amount of information that is revealed by a ciphertext to a given receiver. To capture adversarial correctness based on fine-grained control, verifiable functional encryption[VFE] was defined. Informally speaking, DVFE requires that, for any ciphertext $C$ passing a public verification procedure, there exists a unique message $m$ such that, for any valid function description $f$ and function key $dk_{f}$ passing public verification, the decryption algorithm, when given $C$, $dk_{f}$, and $f$ as inputs, outputs $f(m)$.
We follow the formal definition of VMCFE proposed in [VMCFE] to redefine our DVFE, including eight algorithms as following definition.
Definition 1 (Private-key Decentralized Verifiable Multi-Client Functional Encryption for Inner Product)
A Private-key Decentralized Verifiable Multi-Client Functional Encryption (DVFE) among a set of $n$ clients $\left\{\mathcal{C}_{i}\right\}_{i\in[n]}$ with a $m$-dimensional vector on $\mathcal{M}^{m}$ contains eight algorithms.
  • SetUp($\lambda$) $\to pp$: Takes as input the security parameter $\lambda$ and outputs the public parameters $pp$.
@@ IV-C Verifiable Functional Encryption
  • DKeyGenShare($sk_{i},pk,y_{i},\ell_{y},pp$) $\to\left(dk_{y,i},\pi_{DKeyGenShare,i}\right)$: Takes as input the secret key $sk_{i}$ of a client, public key $pk$, a value $y_{i}$ and a function label $\ell_{y}$. Outputs a functional decryption key share $dk_{i,y}$ and a zero knowledge proof $\pi_{DKeyGenShare,i}$.
@@
  • Decrypt($\left\{C_{\ell,i}\right\}_{i\in[n]},h^{dk_{y}},y$) $\to\left\{\langle x_{1:n,j},y\rangle\right\}_{j\in[m]}$ or $\bot$ : Takes as input the m-dimensions vector ciphertext $\left\{C_{\ell,i}\right\}_{i\in[n]}$ of $n$ client , a functional decryption key $h^{dk_{y}}$ and the function vector $y$. Outputs $\left\{\langle x_{1:n,j},y\rangle\right\}_{j\in[m]}$ or $\bot$.
Correctness. Given any set of message $\left(x_{1},\cdots,x_{n}\right)\in\mathcal{M}^{n\times m}$ and $x_{i},x_{0}\in\mathcal{M}^{m}$ , and any function vector $y\in\mathbb{Z}_{p}^{n}$, We say that an DVFE scheme is correct if and only if (1) holds.
$\Pr\left[\begin{array}[]{l}\begin{aligned} &\texttt{Decrypt}(\{C_{\ell,i}\}_{i\in[n]},h^{dk_{y}},y)\\ &=\left\{\langle x_{1:n,j},y\rangle\right\}_{j\in[m]}\end{aligned}\mid\begin{aligned} &\texttt{SetUp}(\lambda)\rightarrow\mathrm{pp}\\ &\texttt{KeyGen}(\mathrm{pp})\rightarrow\left((\mathrm{sk}_{i},\mathrm{ek}_{i})_{i\in[n]},\;\mathrm{vk}_{\mathrm{CT}},\;\mathrm{vk}_{\mathrm{DK}},\;\mathrm{pk}\right)\\ &\forall i\in[n],\\ &\texttt{Encrypt}\left(\mathrm{ek}_{i},\;x_{i},\;x_{0},\;\ell,\;(\ell_{\mathrm{Enc},j})_{j\in[m]},\;\mathrm{pp}\right)\rightarrow\left(C_{\ell,i},\pi_{\mathrm{Encrypt},i}\right)\\ &\texttt{DKeyGenShare}(\mathrm{sk}_{i},\;\mathrm{pk},\;y_{i},\;\ell_{y},\;\mathrm{pp})\rightarrow\left(dk_{y,i},\pi_{\mathrm{DKeyGenShare},i}\right)\\ &\texttt{VerifyCT}((C_{\ell,i})_{i\in[n]},\;\pi_{Encrypt,i},\;\mathrm{vk}_{\mathrm{CT}})=1\\ &\texttt{VerifyDK}((dk_{y,i})_{i\in[n]},\;\pi_{DKeyGenShare,i},\;\mathrm{vk}_{\mathrm{DK}},\;\mathrm{pp})=1\end{aligned}\end{array}\right]=1.$ (1)
@@ V OUR DECENTRALIZED VERIABLE FUNCTIONAL ENCRYPTION
V-A Construction of DVFE
Figure 1: The workflow and main components of DVFE
@@ V-A Construction of DVFE
  • SetUp($\lambda$) $\to pp$: Takes as input the security parameter $\lambda$ and outputs the public parameters $pp=\left(\mathcal{PG},\mathcal{DG},h_{p},\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}^{\prime}_{1},\mathcal{H},\ell_{D}\right)$, where $\mathcal{PG}:=(\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},p,g,h,e)\overset{\mathdollar}{\leftarrow}\mathrm{\texttt{PGGen}}(\lambda)$ is a pairing group, $g$ the generator of $\mathbb{G}_{1}$, $h$ the generator of $\mathbb{G}_{2}$, $\mathcal{DG}=(\tilde{s},f,\hat{h}_{p},\hat{G},F,\hat{G}^{p})\leftarrow\texttt{GenClassGroup}(\lambda,p)$ is a DDH group with an easy DL subgroup, $\hat{t}\leftarrow\mathcal{D}_{p}$ is sampled, and $h_{p}=\hat{h}_{p}^{\hat{t}}$ is set. $\mathcal{H}_{1}:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}^{2}$, $\mathcal{H}_{2}:\{0,1\}^{*}\rightarrow\mathbb{G}_{2}^{2}$, $\mathcal{H}^{\prime}_{1}:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}$, $\mathcal{H}:\{0,1\}^{*}\rightarrow\mathbb{Z}_{p}$ are four full-domain hash function, and $\ell_{D}:=(\{0,1\}^{*})$ is the initialization label.
  • KeyGen(pppp) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: It is executed by each client $\mathcal{C}_{i}$ to generate own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs a verifiation key for ciphertexts $vk_{CT}$ and functional key share $vk_{Dk}$, and a public key $pk$. It takes $pp$ as input, and a client selects $\hat{k}_{i}=(\hat{k}_{i,1},\hat{k}_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, $t_{i}=(t_{i,1},t_{i,2})\leftarrow\mathcal{D}^{2}_{p}$ and $s_{i}=(s_{i,1},s_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, then computes $T_{i}=\left(h_{p}^{t_{i,b}}\right)_{b\in[2]}$ and
@@
    where $v=\mathcal{H}_{1}(\ell_{D})\in\mathbb{G}_{1}^{2}$.
    For each client $\mathcal{C}_{i}$, encryption key is $ek_{i}=s_{i}$, secret key is $sk_{i}=\left(s_{i},\hat{k}_{i},t_{i}\right)$ , and public key is $pk=\left(T_{i},\ d_{i}\right)_{i\in[n]}$. The verification key for ciphertexts is $vk_{CT}=\left(\mathrm{com}_{i}\right)_{i\in[n]}$, while for functional keys it is $vk_{DK}=\left(T_{i},d_{i},\mathrm{com}_{i}\right)_{i\in[n]}$.
@@
    $\text{PoK}\left\{(s_{i},x_{i}):\begin{aligned} V_{i}=(u^{m})^{s_{i}}\cdot w^{{x}_{i}}\wedge\\ y_{i}=\frac{\langle x_{i},x_{0}\rangle}{\langle x_{i},x_{i}\rangle}\wedge\\ \mathrm{com}_{i}=(v)^{s_{i}}.\end{aligned}\right\}$
    where $w^{{x}_{i}}=\prod_{j\in[m]}(w_{j})^{x_{i,j}}$ and it is detailed in section B. Outputs ciphertext $C_{\ell,i}$ and proof $\pi_{Encrypt,i}$.
@@
    where $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b})\in\mathbb{G}_{2}^{2},b\in[2]$.
    It re-computes $\mathrm{com}_{i}$ and generates a proof $\pi_{DKeyGenShare,i}$ for the relation $\mathcal{R}_{DKeyGenShare,i}$ on input $\left(pk,com,dk_{i,y},\ell_{y}\right)$. The $\pi_{DKeyGenShare,i}$ is described as
@@
    where $K_{\sum,i,b}=\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\in G$, $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b}),b\in[2],v=\mathcal{H}_{1}(\ell_{D})$. It is detailed in section C. Outputs a functional decryption key share $dk_{i,y}$ and $\pi_{DKeyGenShare,i}$.
@@
    where $\hat{v}_{b}=\left(\mathcal{H}_{2}(\ell_{y,b})\right),b\in[2]$. Outputs the functional decryption key $h^{dk_{y}}$.
  • Decrypt($\left\{C_{\ell,i}\right\}_{i\in[n]},h^{dk_{y}},y$) $\to\left(\langle x_{1:n,j},y\rangle\right)_{j\in[m]}$ or $\bot$ : Takes as input $n$ client with m-dimensions vector ciphertext $\left(C_{\ell,i}\right)_{i\in[n]}$, a functional decryption key $h^{dk_{y}}$ and the function vector $y$. It computes $u=(u_{1},u_{2})=\mathcal{H}_{1}(l)$, and $\left(E_{j}\right)_{j\in[m]}=\left(e(w_{j},h)\right)_{j\in[m]}$, where $w_{j}=\mathcal{H}_{1}^{{}^{\prime}}(\ell_{Enc,j})$.
    For $j\in[m]$, it gets
    $E_{j}^{\langle x_{1:n,j},y\rangle}=\frac{\prod_{i\in[n]}e(c_{l,i,j},h^{y_{i}})}{e(u_{1},h^{dk_{y,1}})e(u_{2},h^{dk_{y,2}})}$
    and solves the discrete logarithm in basis $E_{j}$ to return $\langle x_{1:n,j},y\rangle$. It outputs $\left(\langle x_{1:n,j},y\rangle\right)_{j\in[m]}$ or $\bot$.
Correctness.It means that given $pp=\texttt{SetUp}(\lambda)$, $((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk)=\texttt{KeyGen}(pp)$, the ciphertexts with proofs $\{(C_{\ell,i},\pi_{Encrypt,i})\}_{i\in[n]}=\{\texttt{Encrypt}(ek_{i},x_{i},x_{0},\ell,\left(\ell_{Enc,j}\right)_{j\in[m]},pp)\}_{i\in[n]}$, and the decryption key shares $\{(dk_{i,y},\pi_{DKeyGenShare,i})\}_{i\in[n]}=\{\texttt{DKeyGenShare}(sk_{i},pk,y_{i},\ell_{y},pp)\}_{i\in[n]}$,
when $\texttt{VerifyCT}((C_{\ell,i})_{i\in[n]},\;\pi_{Encrypt,i},\;\mathrm{vk}_{\mathrm{CT}})=1$ and $\texttt{VerifyDK}((dk_{y,i})_{i\in[n]},\;\pi_{DKeyGenShare,i},\;\mathrm{vk}_{\mathrm{DK}},\;\mathrm{pp})=1$, we have
@@
$(E_{j})^{\langle x_{1:n,j},y\rangle}$ $=\frac{\prod_{i\in[n]}e(c_{l,i,j},h^{y_{i}})}{e(u_{1},h^{dk_{y,1}})e(u_{2},h^{dk_{y,2}})}$
@@
$=e(w_{j},h)^{\langle x_{1:n,j},y\rangle}$
where $x_{1:n,j}\in\mathbb{R}^{n}$ denote the vector composed of the $j$-th parameter of the local models from each of the $n$ clients, and then solves the discrete logarithm in basis $E_{j}$ to return $\langle x_{1:n,j},y\rangle$. Performing $m$ decryption operations on $\{\left\{c_{\ell,i,j}\right\}_{i\in[n]}\}_{j\in[m]}$ returns the result $\{\langle x_{1:n,j},y\rangle\}_{j\in[m]}$.
@@
$\mathrm{Adv_{DVFE}^{sta-ind}}(t,q_{E},q_{K})\leq\$$q_{E}\mathrm{Adv}^{\mathrm{zk}}_{\prod_{Encrypt}}(t)$
@@
    $\mathrm{Adv_{DVFE}^{sta-ind}}(t,q_{E},q_{K})$ is the best advantage of any PPT adversary running in time $t$ with $q_{E}$ encryption queries and $q_{K}$ key share queries against the static-IND-secure game of the DVFE scheme;
@@
Game $G_{1}$: This game is as $G_{0}$, except that for generating proofs: $\pi_{Encryption}$ and $\pi_{KeyGenShare}$ in the QEncrypt and QDKeyGenShare.
    $\pi_{\mathrm{Encryption}}$ is produced by a simulator $\mathrm{Sim}_{\mathrm{Encrypt}}$ on the input $(\ell,\left\{\ell_{\mathrm{Enc}}\right\}_{j\in[m]},c_{\ell,i},x_{0},\mathrm{com}_{i})$ if $\mathcal{C}_{i}$ is queried in QEncrypt for the input $(i,X^{0},X^{1}\in[0,2^{q}-1]^{n\times m},\ell,\left\{\ell_{\mathrm{Enc}}\right\}_{j\in[m]})$.
@@
    $dk_{i}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{\sum_{t\in\mathcal{HC}}s_{t,1}\cdot y_{t}},\right)_{b\in[2]}$
@@
Game $G_{4}$: We simulate the answers to $\mathcal{H}^{\prime}_{1}$ to any new RO-query of $\mathbb{G}_{1}$ by a truly random element of $w_{j}=g^{\hat{w}_{j}}$, with $\hat{w}_{j}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. $\mathrm{Adv}_{4}=\mathrm{Adv}_{5}$.
@@
For any PPT adversary, the advantage of distinguishing $\{g,g^{r},g^{ar}\mid a\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p},r=RF_{\mathbb{Z}_{p}}(\ell)\}$ and $\{g,g^{u_{1}-u_{2}a},g^{u_{1}a+u_{2}}\mid a,u_{1},u_{2}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\}$ is equal to the advantage of solving the DDH problem.
@@
    $|\operatorname{Pr}[b=0]\left(\operatorname{Pr}\left[G_{5.q.t}^{\star}\middle|b=0,E\right]\cdot\operatorname{Pr}[E]+\frac{\operatorname{Pr}[\neg E]}{2}\right)$
    $-\operatorname{Pr}[b=1]\left(\operatorname{Pr}\left[G_{5.q.t}^{\star}\middle|b=1,E\right]\cdot\operatorname{Pr}[E]+\frac{\operatorname{Pr}[\neg E]}{2}\right)|$
    $=$ $\operatorname{Pr}[E]\cdot|\operatorname{Pr}[b=0]\cdot\operatorname{Pr}\left[G_{5.q\cdot t}^{\star}\middle|b=0,E\right]$
    $-\operatorname{Pr}[b=1]\cdot\operatorname{Pr}\left[G_{5.q\cdot t}^{\star}\middle|b=1,E\right]|$

### v4

# VFEFL: Privacy-preserving federated learning against malicious clients via verifiable functional encryption

@@ Abstract
Federated learning is a promising distributed learning paradigm that enables collaborative model training without exposing local client data, thereby protecting data privacy. However, it also brings new threats and challenges. The advancement of model inversion attacks has rendered the plaintext transmission of local models insecure, while the distributed nature of federated learning makes it particularly vulnerable to attacks raised by malicious clients. To protect data privacy and prevent malicious client attacks, this paper proposes a privacy-preserving Federated Learning framework based on Verifiable Functional Encryption (VFEFL), without a non-colluding dual-server assumption or additional trusted third-party. Specifically, we propose a novel Cross-Ciphertext Decentralized Verifiable Functional Encryption (CC-DVFE) scheme that enables the verification of specific relationships over multi-dimensional ciphertexts. This scheme is formally treated, in terms of definition, security model and security proof. Furthermore, based on the proposed CC-DVFE scheme, we design a privacy-preserving federated learning framework that incorporates a novel robust aggregation rule to detect malicious clients, enabling the effective training of high-accuracy models under adversarial settings. Finally, we provide the formal analysis and empirical evaluation of VFEFL. The results demonstrate that our approach achieves the desired privacy protection, robustness, verifiability and fidelity, while eliminating the reliance on non-colluding dual-server assumption or trusted third parties required by most existing methods.
@@ III-A System Model
Note: The superscripts of $W_{i}^{t},\vec{x}_{i}^{t},D_{i}$ denote interactions and the subscripts denote client identifiers.
@@ III-C Security Model
where $\stackrel{{\scriptstyle c}}{{\approx}}$ denotes computational indistinguishability, $\lambda$ represents the security parameter, and $\hat{x}$ includes all inputs.
@@ IV-A Groups
Prime Order Group. We use a prime-order group GGen, a probabilistic polynomial-time (PPT) algorithm that, given a security parameter $\lambda$, outputs a description $\mathcal{G}=(\mathbb{G},p,g)$, where $\mathbb{G}$ is an additive cyclic group of order $p$, $p$ is a $2\lambda$-bit prime, and $g$ is the generator of $\mathbb{G}$. For any $a\in\mathbb{Z}_{p}$, given $g^{a}$, it is computationally hard to determine $a$, a problem known as the discrete logarithm problem. For any $\vec{a}=\{a_{1},\cdots,a_{n}\}\in\mathbb{Z}^{n}_{p}$, $\vec{g}=\{g_{1},\cdots,g_{n}\}\in\mathbb{G}^{n}$, we denote $\vec{g}^{\vec{a}}=\prod_{i\in[n]}g_{i}^{a_{i}}$.
@@ IV-B Zero Knowledge Proof
IV-C Cross-Ciphertext Decentralized Verifiable Functional Encryption
Verifiable functional encryption (VFE) [VFE] was introduced to capture ciphertexts and decryption keys’ correctness under fine-grained control. Informally, VFE requires that for any ciphertext $C$ that passes a public verification procedure, there exists a unique message $m$ such that, for any valid function description $f$ and any corresponding function key $dk_{f}$ that also pass public verification, the decryption algorithm outputs $f(m)$ when given $C$, $dk_{f}$, and $f$ as inputs.
Nguyen et al.[VMCFE] proposed a decentralized multi-client verifiable functional encryption scheme. The scheme focuses on verifying properties of individual ciphertexts. In our work, we propose a verifiable functional encryption scheme that enables cross-ciphertext verification, allowing the validation of specific relations among multiple ciphertexts in various application scenarios. Next, we give the formal definition of our Cross-Ciphertext Decentralized Verifiable multi-client Functional Encryption for inner product (CC-DVFE).
Definition 1
A Cross-Ciphertext Decentralized Verifiable Multi-Client Functional Encryption for Inner Product among a set of $n$ clients $\left(\mathcal{C}_{i}\right)_{i\in[n]}$ with a $m$-dimensional vector on $\mathcal{M}^{m}$ contains eight algorithms.
[…]
@@ IV-C Cross-Ciphertext Decentralized Verifiable Functional Encryption
[…]
Correctness. Given any set of message $\left(\vec{x}_{1},\cdots,\vec{x}_{n}\right)\in\mathcal{M}^{n\times m}$ and the auxiliary information auxaux and any function vector $\vec{y}\in\mathbb{Z}_{p}^{n}$, We say that an CC-DVFE scheme is correct if and only if following equation holds.
$\begin{split}\Pr\left[\begin{array}[]{l|l}&\texttt{SetUp}(\lambda)\rightarrow\mathrm{pp}\\ &\texttt{KeyGen}(pp)\rightarrow((ek_{i},sk_{i})_{i\in[n]},\\ &\quad vk_{CT},vk_{DK},pk),\\ \texttt{DKeyComb}(&\forall i\in[n],\ \texttt{Encrypt}(ek_{i},\vec{x}_{i},\\ \quad({dk}_{y,i})_{i\in[n]},\ell_{y},&\quad aux,\ell,(\ell_{Enc,j})_{j\in[m]},pp)\\ \quad pk)\to dk&\quad\rightarrow(C_{\ell,i},\pi_{CT,i})\\ \texttt{Decrypt}((C_{\ell,i})_{i\in[n]},&\forall i\in[n],\ \texttt{DKeyGenShare}(\\ \quad{dk},\vec{y})\to&\quad sk_{i},pk,y_{i},\ell_{y},pp)\\ \quad\left(\langle\vec{x}^{\top},\vec{y}\rangle\right)_{j\in[m]}&\quad\rightarrow(dk_{y,i},\pi_{\mathrm{DK},i})\\ &\texttt{VerifyCT}((C_{\ell,i})_{i\in[n]},\pi_{CT},\\ &\quad vk_{CT})\to 1\\ &\texttt{VerifyDK}(({dk}_{y,i})_{i\in[n]},\pi_{DK},\\ &\quad vk_{DK},pp)\to 1\\ \end{array}\right]=1.\end{split}$
@@ V OUR DECENTRALIZED VERIABLE FUNCTIONAL ENCRYPTION
V-A Construction of CC-DVFE
Figure 1: The workflow and main components of CC-DVFE
@@ V-A Construction of CC-DVFE
  • SetUp($\lambda$) $\to pp$: Takes as input the security parameter $\lambda$ and outputs the public parameters $pp=(\mathcal{PG},\mathcal{DG},h_{p},\mathcal{H}_{1},\mathcal{H}_{2},$ $\mathcal{H}^{\prime}_{1},\mathcal{H},\ell_{D})$, where $\mathrm{\texttt{PGGen}}(\lambda)\to\mathcal{PG}=(\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},p,g,h,e)$ is a pairing group, $g$ the generator of $\mathbb{G}_{1}$, $h$ the generator of $\mathbb{G}_{2}$, GenClassGroup $(\lambda,p)\to\mathcal{DG}=(\tilde{s},f,\hat{h}_{p},$ $\hat{G},F,\hat{G}^{p})$ is a DDH group with an easy DL subgroup, $\hat{t}\leftarrow\mathcal{D}_{p}$ is sampled, and $h_{p}=\hat{h}_{p}^{\hat{t}}$ is set. Four full-domain hash functions are defined as follows:
    $\mathcal{H}_{1}$ $:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}^{2},$ $\mathcal{H}_{2}$ $:\{0,1\}^{*}\rightarrow\mathbb{G}_{2}^{2},$
    $\mathcal{H}^{\prime}_{1}$ $:\{0,1\}^{*}\rightarrow\mathbb{G}_{1},$ $\mathcal{H}$ $:\{0,1\}^{*}\rightarrow\mathbb{Z}_{p}.$
    And $\ell_{D}=(\{0,1\}^{*})$ is the initialization label.
  • KeyGen(pppp) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: It is executed by each client $\mathcal{C}_{i}$ to generate own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs a verifiation key for ciphertexts $vk_{CT}$ and functional key share $vk_{Dk}$, and a public key $pk$. A client selects $\hat{k}_{i}=(\hat{k}_{i,1},\hat{k}_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, $t_{i}=(t_{i,1},t_{i,2})\overset{\mathdollar}{\leftarrow}\mathcal{D}^{2}_{p}$ and $s_{i}=(s_{i,1},s_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, then computes $T_{i}=\left(h_{p}^{t_{i,b}}\right)_{b\in[2]}$ and
@@
    where $v=\mathcal{H}_{1}(\ell_{D})\in\mathbb{G}_{1}^{2}$. For each client $\mathcal{C}_{i}$, encryption key is $ek_{i}=s_{i}$, secret key is $sk_{i}=(s_{i},\hat{k}_{i},t_{i})$ , and public key is $pk=\left(T_{i},\ d_{i}\right)_{i\in[n]}$. The verification key for ciphertexts is $vk_{CT}=\left(\mathrm{com}_{i}\right)_{i\in[n]}$, while for functional keys it is $vk_{DK}=\left(T_{i},d_{i},\mathrm{com}_{i}\right)_{i\in[n]}$.
@@
    $\text{PoK}\left\{(s_{i},\vec{x}_{i}):\begin{aligned} V_{i}=(u^{m})^{s_{i}}\cdot\vec{w}^{\vec{x}_{i}}\wedge\\ y_{i}=\frac{\langle\vec{x}_{i},\vec{x}_{0}\rangle}{\langle\vec{x}_{i},\vec{x}_{i}\rangle}\wedge\\ \mathrm{com}_{i}=(v)^{s_{i}}.\end{aligned}\right\}$
    where $\vec{w}^{\vec{x}_{i}}=\prod_{j\in[m]}(w_{j})^{x_{i,j}}$. Outputs ciphertext $C_{\ell,i}$ and proof $\pi_{CT,i}$.
@@
    where $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b})\in\mathbb{G}_{2}^{2},b\in[2]$. It re-computes $\mathrm{com}_{i}$ and generates a proof $\pi_{DK,i}$ for the relation $\mathcal{R}_{DKeyGenShare,i}$ on input $\left(pk,com,{dk}_{y,i},\ell_{y}\right)$. The $\pi_{DK,i}$ is described as
@@
    where $K_{\sum,i,b}=\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\in G$, $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b}),b\in[2],v=\mathcal{H}_{1}(\ell_{D})$. Outputs a functional decryption key share ${dk}_{y,i}$ and $\pi_{DK,i}$.
@@
    where $\hat{v}_{b}=\left(\mathcal{H}_{2}(\ell_{y,b})\right),b\in[2]$. Outputs the functional decryption key $dk=h^{dk_{y}}$.
[…]
    $E_{j}^{\langle\vec{x}^{\top},\vec{y}\rangle}=\frac{\prod_{i\in[n]}e(c_{l,i,j},h^{y_{i}})}{e(u_{1},h^{dk_{y,1}})e(u_{2},h^{dk_{y,2}})}$
    and solves the discrete logarithm in basis $E_{j}$ to return $\langle\vec{x}^{\top},\vec{y}\rangle$, where $\vec{x}^{\top}=(x_{1,j},x_{2,j},\dots,x_{n,j})\in\mathcal{M}^{n}$ denotes the vector consisting of the $j$-th component of each of $n$ clients. It outputs $(\langle\vec{x}^{\top},\vec{y}\rangle)_{j\in[m]}$ or $\bot$.
Correctness.We have
@@
$(E_{j})^{\langle\vec{x}^{\top},y\rangle}$ $=\frac{\prod_{i\in[n]}e(c_{l,i,j},h^{y_{i}})}{e(u_{1},h^{dk_{y,1}})e(u_{2},h^{dk_{y,2}})}$
@@
$=e(w_{j},h)^{\langle\vec{x}^{\top},y\rangle}$
where $\vec{x}^{\top}=(x_{1,j},x_{2,j},\dots,x_{n,j})\in\mathcal{M}^{n}$ denote the vector composed of the $j$-th parameter of the local models from each of the $n$ clients, and then solves the discrete logarithm in basis $E_{j}$ to return $\langle\vec{x}^{\top},y\rangle$. Performing $m$ decryption operations on $\{\left\{c_{\ell,i,j}\right\}_{i\in[n]}\}_{j\in[m]}$ returns the result $\{\langle\vec{x}^{\top},y\rangle\}_{j\in[m]}$.
@@
$\mathrm{Adv_{CC-DVFE}^{sta-ind}}(t,q_{E},q_{K})\leq\$$q_{E}\mathrm{Adv}^{\mathrm{zk}}_{\prod_{Encrypt}}(t)$
@@
    $\mathrm{Adv_{CC-DVFE}^{sta-ind}}(t,q_{E},q_{K})$ is the best advantage of any PPT adversary running in time $t$ with $q_{E}$ encryption queries and $q_{K}$ key share queries against the static-IND-secure game of the CC-DVFE scheme;
@@
Game $G_{1}$: This game is as $G_{0}$, except that for generating proofs: $\pi_{CT}$ and $\pi_{DK}$ in the QEncrypt and QDKeyGenShare.
    $\pi_{\mathrm{CT}}$ is produced by a simulator $\mathrm{Sim}_{\mathrm{Encrypt}}$ on the input $(\ell,\left\{\ell_{\mathrm{Enc}}\right\}_{j\in[m]},C_{\ell,i},\vec{x}_{0},\mathrm{com}_{i})$ if $\mathcal{C}_{i}$ is queried in QEncrypt for the input $(i,X^{0},X^{1}\in[0,2^{q}-1]^{n\times m},\ell,\left\{\ell_{\mathrm{Enc}}\right\}_{j\in[m]})$.
@@
    $dk_{i}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{\sum_{t\in\mathcal{HC}}s_{t,1}\cdot y_{t}}\right)_{b\in[2]}$
@@
Game $G_{4}$: We simulate the answers to $\mathcal{H}^{\prime}_{1}$ to any new RO-query of $\mathbb{G}_{1}$ by a truly random element of $w_{j}=g^{\hat{w}_{j}}$, with $\hat{w}_{j}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. $\mathrm{Adv}_{3}=\mathrm{Adv}_{4}$.
@@
For any PPT adversary, let $\mathrm{Adv}(t)$ denote its advantage in distinguishing the distributions $\mathcal{D}_{0}=\{g,g^{r},g^{ar}\mid a\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p},\ r=RF_{\mathbb{Z}_{p}}(\ell)\}$ and $\mathcal{D}_{1}=\{g,g^{u_{1}-u_{2}a},g^{u_{1}a+u_{2}}\mid a,u_{1},u_{2}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\}$. Let $\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$ denote the advantage in solving the DDH problem in $\mathbb{G}_{1}$. Then it holds that $\mathrm{Adv}(t)\leq 2\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$.
@@
    $|\operatorname{Pr}\left[G_{5.q.t}^{\star}\middle|b=0,E\right]\cdot\operatorname{Pr}[E]+\frac{\operatorname{Pr}[\neg E]}{2}$
    $-\operatorname{Pr}\left[G_{5.q.t}^{\star}\middle|b=1,E\right]\cdot\operatorname{Pr}[E]+\frac{\operatorname{Pr}[\neg E]}{2}|$
    $=$ $\operatorname{Pr}[E]\cdot|\operatorname{Pr}\left[G_{5.q\cdot t}^{\star}\middle|b=0,E\right]-\operatorname{Pr}\left[G_{5.q\cdot t}^{\star}\middle|b=1,E\right]|$